OpenStudy (anonymous):

True or false: The equation sec2x - 2 = 0 has 4 solutions on the interval [0, 2pi). arcsec(2)/2= 0.523598775 0.523598775+pi=3.6651914285897932384626433832795 Did i miss anything?

OpenStudy (anonymous):

also 2pi-pi/6= 5.7595865315812876038481795360124 Is there a fourth?

OpenStudy (anonymous):

5.7595865315812876038481795360124-pi is one but is there a fifth

OpenStudy (anonymous):

\[\sec 2x=2\]\[\frac{ 1 }{ \cos 2x }=2\] \[\cos 2x=\frac{ 1 }{ 2 }=\cos \frac{ \pi }{ 3 },\cos \left( 2 \pi-\frac{ \pi }{ 3 } \right),\cos \left( 2 \pi+\frac{ \pi }{ 3 } \right), \cos \left( 2 \pi+2 \pi-\frac{ \pi }{ 3 } \right)\] \[2 x=\frac{ \pi }{ 3 },\frac{ 5 \pi }{ 3 },\frac{ 7 \pi }{ 3 },\frac{ 11 \pi }{ 3 }\] x=?

OpenStudy (anonymous):

Exactlly what i had, thanks :)